OpenStudy (anonymous):

Solve x(t) for 5x’(t)+3x(t)=6, x(0)=1, where x’(t) is the derivative of x(t) with respect to t.

OpenStudy (anonymous):

put it in standard form and use integrating factor

OpenStudy (anonymous):

or wolframalpha

OpenStudy (anonymous):

im confuse

sam (.sam.):

Try these http://www.khanacademy.org/#differential-equations

OpenStudy (anonymous):

x'(t) + (3/5)x(t)= 6/5 x'(t) + (3/5)x(t) - 6/5= 0 (1/2)x(t) + (3/5)x(t) - 6/5= 0 (1/2)x(t) + (3/5)x(t) - 6/5= 0 ??

OpenStudy (ash2326):

plpeter673 that's not correct,which grade you are in high school?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

who's asking

OpenStudy (ash2326):

I'll help you according to your grade , can be solved in several ways

OpenStudy (anonymous):

yes, hs

OpenStudy (ash2326):

okay, Do you know what's integrating factor?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

we have \[5x'(t)+3x(t)=6\]Let's write this as \[ 5\frac{dx}{dt}+3 x=6\] where x= x(t) \[ 5 \frac{dx}{dt}= 6-3x\] This can be written as \[ \frac{dx}{6-3x}= \frac{dt}{5}\] now integrate both sides We know integration of \(\frac{1}{a-x}\)= \(-\ln(a-x)\) here ln= log to the base e We get \[ -\frac{1}{3} \times \ln(6-3x) = \frac{t}{5}+C\] C= constant of integration We get \[ \ln {6-3x} = -\frac{3t}{5} +C\] Take antilog both sides \[ 6-3x= e^{-\frac{3t}{5}}\times C\] we get x(t) \[x(t)= \frac{6-C\times e^{-\frac{3t}{5}}}{3} \] we are given at t=0 x(t)=1 so \[1=\frac{6-C\times1}{3} \] \[3=6-C\] we get \[C= 3\] so \[x(t)= \frac{6-3\times e^{-\frac{3t}{5}}}{3} \]

OpenStudy (ash2326):

x(t) can also be written as \[x(t)=2-e^{-\frac{3t}{5}}\]

OpenStudy (anonymous):

cool. thk

OpenStudy (ash2326):

welcome :) Did you get this?

OpenStudy (anonymous):

yes